Suppose f varies directly as the square of m, and inversely as j. Also, if f=36 when m=2 and j=10. Find f, if m=4 and j=6.
Answer by Theo(13342)

f varies directly as the square of m and inversely as j.

if f = 36 when m = 2 and j = 10, find the value of f when m = 4 and j = 6.

the formula is f = k * m^2 / j.

when f = 36 and m = 2 and j = 10, the formula becomes:

36 = k * 2^2 / 10.

solve for k to get:

k = 36 * 10 / 2^2 = 90.

k is the constant of variation.
it does not change it's value in the equation, while all the other variables can change their value.

when m = 4 and j = 6, the formula becomes:

f = k * m^2 / j becomes:

f = 90 * 4^2 / 6.

solve for f to get:

f = 240.

that's your solution.